# Improving the fact-checking performance of language models by relying on their entailment ability

### v1

@@ 4.3 Observations from TBE-3:
Table 5: Performance of Prompting methods across IBE-1, IBE-2, and IBE-3 settings for LIAR-RAW and RAW-FC datasets.
4.4 Observations from IBEs’:
We reported the performance of various models in IBEs in Table 11. For LIAR-RAW dataset, Mistral achieved the highest macro-F1 (0.22) in IBE-1. While in IBE-2, Llama achieved the highest performance ($MF1:\textbf{0.22}$), Mistral, Llama, Qwen and Falcon ($MF1:\textbf{0.21}$) gave the highest performance in IBE-3. In IBE-4, Mistral ($MF1:\textbf{0.14}$) and Falcon ($MF1:\textbf{0.14}$) got the highest performance. None of the macro-F1 could surpass the baseline performances. In contrast, for RAW-FC dataset, Llama ($MF1:\textbf{0.62}$) achieved the highest performace in IBE-2, surpassing the highest baseline performance ($MF1:$ 0.61). While Qwen attained the highest performance in IBE-1 ($MF1:$ 0.59), surpassing the baselines HiSS ($MF1:$ 0.53), FactLLaMa ($MF1:$ 0.55) and RAFTS ($MF1:$ 0.57), it is still behind the performance of L-Defense ($MF1:$ 0.61). Similarly, Qwen got the highest performance in IBE-3 ($MF1:$ 0.52). In IBE-4, Mistral and Qwen achieved the highest macro-F1 of 0.43, which is far behind all baselines. We have reported some additional observations in the Appendix (section D.4)
4.5 Comparing the performances across TBEs and IBEs:
In this section, we have reported our findings by comparing the best-performing models across all TBEs and IBEs. The comparison is also illustrated in Figure 22, Figure 23 and Figure 24 in the Appendix D.5. For the LIAR-RAW dataset, we found that none of the models from any IBEs, TBE-1 and TBE-2 ($MF1$: 0.14 - 0.32) could surpass the highest F1 achieved in baselines ($MF1$: 0.42). However, we observed that TBE-1 and TBE-2 ($MF1$: 0.30 - 0.32, $\sim\textbf{45.45}\%\uparrow$) outperformed IBEs ($MF1$: 0.22) in terms of the highest reported macro-F1. Further, we also found that many models from TBE-3 ($MF1$: 0.44 - 0.54, $\sim\textbf{68.75}\%\uparrow$) outperform all of the models of IBEs ($MF1$: 0.14 - 0.22), TBE-1 ($MF1$: 0.30) and TBE-2 ($MF1$: 0.32). For the RAW-FC dataset, we observed a similar comparative performance. However, here, the highest reported macro-F1 by IBE-2 (0.62) is comparable to the best-performing baseline ($MF1$: 0.61). Similar to the LIAR-RAW dataset, here also we observed that (i) many TBE-1 and TBE-2 models ($MF1$: 0.62 - 0.71, $\sim\textbf{14.52}\%\uparrow$) outperform IBEs (0.62) in terms of macro-F1, and (ii) many TBE-3 models ($MF1$: 0.64 - 0.88, $\sim\textbf{23.94}\%\uparrow$) outperform all of the models of IBEs ($MF1$: 0.43 - 0.62), TBE-1 ($MF1$: 0.65 - 0.66) and TBE-2 ($MF1$: 0.62 - 0.71). The whole comparison points to the fact that training the LLMs with VLLM-generated entailed-justifications can improve the performance of veracity prediction. We have reported some additional observations in the section D.5 of the appendix due to space constraints. Further, we have also presented the confusion metrics of best-performing TBE and IBE models for LIAR-RAW and RAW-FC datasets in Figure 37 and 38 respectively.
4.6 Insights from evaluation of explanations:
In this section, we reported our findings from evaluating model explanations. We obtained them by concatenating the supporting and refuting justifications. The details of the metrics and evaluation scores are provided in Appendix C.6.4 and Appendix D.6, respectively. We observed that lexical overlap measures like ROUGE ($R_{1}$, $R_{2}$, $R_{L}$) provided mixed signals. For instance, while Falcon generated explanations showed higher unigram-based overlaps ($R_{1}$: 0.23 for LIAR-RAW and $R_{1}$: 0.40 for RAW-FC), Mistral generated explanations got the highest overlap in longest common subsequences ($R_{L}$: 0.14 for LIAR-RAW and $R_{L}$: 0.18 for RAW-FC). However, all models scored consistently low on the BLEU score and the BERT score. On the other hand, subjective evaluations indicated that Llama-generated explanations were rated highest by the majority of language models for dimensions like informativeness, readability, objectivity, and logicality for both datasets. These high subjective ratings seem to be correlated with better veracity prediction performance, as RoBERTa and XLNet performed well upon taking Llama explanations. A more detailed discussion on the observations of ‘evaluation of explanations’ is presented in Appendix D.6.
@@ 5 Conclusion:
  • The role of VLLM entailed justification as a second step in TBE-3 is justified in the ablation study (see Appendix E) where we observed that removing supporting and refuting justification adversarially affected the scores.
  • In the linguistic analysis of model explanations, we found that LLM could attend to supporting and refuting keywords and factual pieces of information for samples labelled with ‘true’ and ‘false’ veracity samples. However, for samples with other veracity labels LLM attention seem to be scattered. A detailed discussion is presented in Appendix F due to space constraints.
@@ Appendix A Related works:
The NLP community has long been focusing on the core and peripheral tasks of automated fact-checking. Several datasets and methodologies have been proposed in the past, spanning different domains, languages, and task frameworks. In the following, we have reported the past literature for (i) monolingual fact-checking, (ii) multi-lingual and multi-modal fact-checking, (iii) explainable fact-checking, and (iv) associated tasks in this area. Apart from that, we have also reported how our work fills the current research gap.
@@ Appendix B Additional details on datasets:
Figure 2: Bar plot showing number of claims v/s number of evidences for LIAR-RAW dataset
@@
Table 6: Example entry from the LIAR-RAW dataset. It contains five key fields: (i) an identifier ‘Event_id’, (ii) the ‘claim’ to be fact-checked, (iii) the ground-truth ‘Label’, (iv) an explanation (‘Explain’) that clarifies its ground truth, and (v) the evidence sentences (E1–E4) extracted from the underlying fact-checking article.
@@
Table 7: Example entry from the RAW-FC dataset. It contains five key fields: (i) an identifier ‘Event_id’, (ii) the ‘claim’ to be fact-checked, (iii) the ground-truth ‘Label’, (iv) an explanation (‘Explain’) that clarifies its ground truth, and (v) the evidence sentences (E1–E4) extracted from the underlying fact-checking article.
Figure 3: Bar plot showing number of claims v/s number of evidences for RAW-FC dataset
@@ C.4 IBE-4: Prompting based on entailment:
In the last inference-based experiment, we tried to answer "Can prompting with VLLM generated entailed-justifications enhance language models’ ability to predict claim veracity?" To conduct this experiment, we have followed a three-step approach similar to TBE-3 as described in Section 3.1.3. While the initial two steps, i.e. (i) to classify the evidence sentences as supporting or refuting a given claim and (ii) generating the justifications based on classified evidence sentences, are exactly the same, in the last step, instead of training, we prompted the VLLMs to generate the veracity. In past, researchers have prompted the VLLMs to find entailment for tasks like claim matching Choi and Ferrara (2024b) and counterfactual generation Dai et al. (2022). However, to the best of our knowledge, we are the first to (i) apply it to classify each evidence into supporting or refuting categories, and (ii) generate entailed justifications and use them for fact verification. The detailed approach is illustrated in Fig. 1 (c). Some of the prompt samples are show cased in the Figure 12, Figure 13, Figure 14, Figure 15 and 16.
@@ C.6 Experimental set-up:
Figure 17: A schematic diagram explaining the TBE-3 pipeline which breaks down into three stages. (a) Entailment: Given a claim and its corresponding evidences, a VLLM first distinguishes evidences into supporting/ refuting evidence via entailment. (b) Evidence consolidation: Then the same VLLM consolidates these two groups into concise supporting and refuting justification. (c) Veracity prediction: Using the claim and both justifications, an LLM is trained to predict veracity.
@@ C.6.4 Evaluation metrics:
Since we are working in a multi-class classification framework for veracity prediction, we used standard metrics such as macro-precision ($MP$), macro-recall ($MR$), and macro-f1 ($MF$) to evaluate our models. To evaluate model explainability, we first concatenated the supporting and refuting entailed justifications (generated as part of TBE-3) and considered them as model explanations. The evaluation was done with two types of strategies: (i) checking lexical overlap and semantic matching, and (ii) doing subjective evaluation by VLLMs. To check lexical overlap, we used several standard evaluation metrics such as ROUGE-1 ($R_{1}$), ROUGE-2 ($R_{2}$), ROUGE-L ($R_{L}$) Lin (2004) and BLEU Papineni et al. (2002). While $R_{1}$ and $R_{2}$ measure the overlap of unigrams and bigrams between predicted and gold explanations, $R_{L}$ measures the longest common subsequence. BLEU, on the other hand, measures the precision of matching n-grams between predicted and gold explanations. To measure the semantic matching between predicted and gold explanations, we deployed BERT score Zhang et al. (2019). It generates contextual embeddings of predicted and gold explanations and calculates cosine similarity between them. To do the subjective evaluation, we prompted the considered VLLMs (a.k.a evaluating VLLMs) to asses the model explanations generated by each VLLM (generating VLLM). The VLLMs were asked to asses across five dimensions (i) informativeness, (ii) logicality, (iii) objectivity, (iv) readability and (v) accuracy Zheng et al. (2025). The prompt template used for the assessment is presented in Figure 25.
@@ Appendix D Additional Results and Discussion:
D.1 Additional observations from TBE-1:
@@
  • We found that Mistral trained with LoRA adapter resulted in the highest macro-precision (0.44) for the LIAR-RAW dataset. While it surpasses the performance of baselines such as FactLLaMA ($MP:$ 0.32, $\textbf{37.50}\%\uparrow$) and L-Defense ($MP:$ 0.31, $\sim\textbf{41.94}\%\uparrow$), it is still behind HiSS ($MP:$ 0.46) and RAFTS ($MP:$ 0.47). For RAWFC dataset, however, many models, such as Mistral trained with LoRA adapter ($MP:$ 0.69, $\sim\textbf{11.29}\%\uparrow$), Llama trained with LoRA ($MP:$ 0.68, $\sim\textbf{9.68}\%\uparrow$) and LoRA+ adapters ($MP:$ 0.67, $\sim\textbf{8.06}\%\uparrow$), Qwen trained with LoRA ($MP:$ 0.67, $\sim\textbf{8.06}\%\uparrow$) and LoRA+ ($MP:$ 0.70, $\sim\textbf{12.90}\%\uparrow$) adapters, outperformed the best performance reported by the baselines ($MP:$ 0.62). Qwen trained with LoRA+ adapter achieved the highest overall performance ($MP:$ 0.70).
  • We observed that Llama trained with LoRA adapter resulted in the highest macro-recall (0.30) for the LIAR-RAW dataset. It falls behind the macro-recall reported by all baselines HiSS ($MR:$ 0.31), FactLLaMa ($MR:$ 0.32), L-Defense ($MR:$ 0.31), and RAFTS ($MR:$ 0.37). However, for the RAW-FC dataset, we observed that many models, such as Mistral trained with LoRA adapter ($MR:$ 0.65, $\sim\textbf{6.56}\%\uparrow$), Llama trained with LoRA ($MR:$ 0.64, $\sim\textbf{4.92}\%\uparrow$) and LoRA+ adapters ($MR:$ 0.64, $\sim\textbf{4.92}\%\uparrow$), Qwen trained with LoRA adapter ($MR:$ 0.66, $\sim\textbf{8.20}\%\uparrow$) outperformed the best reported macro-recall score by the baselines (0.61). Qwen trained with LoRA adapter achieved the highest overall performance ($MR:$ 0.66).
D.2 Additional observations from TBE-2:
@@
  • We found that Mistral trained (with Mistral understandings) with the LoRA adapter resulted in the highest macro-precision (0.36) for the LIAR-RAW dataset. While it outperforms the baselines FactLLaMA ($MP:$ 0.32) and L-Defense ($MP:$ 0.31) models, it is still behind HiSS ($MP:$ 0.46) and RAFTS ($MP:$ 0.47). However, for the RAW-FC dataset, we observed that many models, such as XLNet fine-tuned with Llama understandings ($MP:$ 0.63, $\sim\textbf{1.61}\%\uparrow$), Mistral trained (with Mistral understandings) with LoRA+ ($MP:\textbf{0.66}$, $\sim\textbf{6.45}\%\uparrow$) and Llama trained (with Llama understandings) with LoRA+ ($MP:\textbf{0.73}$, $\sim\textbf{17.74}\%\uparrow$) adapters outperformed the best reported macro-precision by the baselines ($MP:$ 0.62). Llama trained (with Llama understandings) with LoRA+ adapter resulted in the overall highest macro-precision ($MP:$ 0.73).
  • We found that Mistral trained (with Mistral understandings) with LoRA adapter gave us the highest macro-recall (0.32) for LIAR-RAW dataset. While this performance is comparable to the performance of baselines HiSS ($MR:\textbf{0.31}$), FactLLaMA ($MR:\textbf{0.32}$), and L-Defense ($MR:\textbf{0.32}$), it is still behind RAFTS ($MR:\textbf{0.37}$). However, for the RAW-FC dataset, we observed that many models, such as Llama trained (with Llama understandings) with LoRA+ adapter ($MR:$ 0.71, $\sim\textbf{16.39}\%\uparrow$) and XLNet fine-tuned with Llama understandings ($MR:$ 0.63, $\sim\textbf{3.27}\%\uparrow$) outperformed the best reported macro-recall score by the baselines ($MR:\textbf{0.61}$). Llama trained (with Llama understandings) with LoRA+ adapter achieved the highest macro-recall overall ($MR:$ 0.71 $\sim\textbf{16.39}\%\uparrow$).
D.3 Additional observations from TBE-3:
  • We found that XLNet fine-tuned with Llama entailed justifications gave the highest macro-precision (0.55) for the LIAR-RAW dataset. Many models, such as RoBERTa fine-tuned with Mistral ($MP:\textbf{0.48}$, $\sim\textbf{2.12}\%\uparrow$), Llama ($MP:\textbf{0.53}$, $\sim\textbf{12.76}\%\uparrow$), Gemma ($MP:\textbf{0.49}$, $\sim\textbf{4.25}\%\uparrow$), Qwen ($MP:\textbf{0.48}$, $\sim\textbf{2.12}\%\uparrow$), and Falcon ($MP:\textbf{0.49}$, $\sim\textbf{4.25}\%\uparrow$) entailed justifications, XLNet fine-tuned with Mistral ($MP:\textbf{0.49}$, $\sim\textbf{4.25}\%\uparrow$), Llama ($MP:\textbf{0.55}$, $\sim\textbf{17.02}\%\uparrow$), and Qwen ($MP:\textbf{0.50}$, $\sim\textbf{6.38}\%\uparrow$) entailed justifications and Llama trained (with Llama entailed justifications) with LoRA+ adapter ($MP:\textbf{0.49}$, $\sim\textbf{4.25}\%\uparrow$) surpassed the best reported macro-precision score of baselines ($MP:\textbf{0.47}$). Similarly, for the RAW-FC dataset, we observed that many models, such as such as RoBERTa fine-tuned with Mistral ($MP:\textbf{0.83}$, $\sim\textbf{33.87}\%\uparrow$), Llama ($MP:\textbf{0.88}$, $\sim\textbf{41.93}\%\uparrow$), Qwen ($MP:\textbf{0.73}$, $\sim\textbf{17.74}\%\uparrow$), and Falcon ($MP:\textbf{0.65}$, $\sim\textbf{4.83}\%\uparrow$) entailed justifications, XLNet fine-tuned with Mistral ($MP:\textbf{0.83}$, $\sim\textbf{33.87}\%\uparrow$), Llama ($MP:\textbf{0.88}$, $\sim\textbf{41.93}\%\uparrow$), Qwen ($MP:\textbf{0.70}$, $\sim\textbf{12.90}\%\uparrow$), and Falcon ($MP:\textbf{0.76}$, $\sim\textbf{22.58}\%\uparrow$) entailed justifications, Mistral trained (with Mistral entailed justifications) with LoRA ($MP:\textbf{0.69}$, $\sim\textbf{11.29}\%\uparrow$), Llama trained (with Llama entailed justifications) with LoRA ($MP:\textbf{0.63}$, $\sim\textbf{1.61}\%\uparrow$) and LoRA+ adapters ($MP:\textbf{0.84}$, $\sim\textbf{35.48}\%\uparrow$), respectively, surpassed the best macro-precision reported by the baselines ($MP:\textbf{0.61}$). RoBERTa and XLNet fine-tuned with Llama entailed justifications gave the highest macro-precision ($MP:\textbf{0.88}$).

### v2

@@ 4.1 Observations for veracity prediction:
Table 5: Performance of Prompting methods across IBE-1 to IBE-4 settings for LIAR-RAW and RAW-FC datasets in terms of MF1. Green and Blue indicate best and second-best performance, respectively.
4.2 Observations from model explainability:
In this section, we have reported our observations from evaluating model explanations. The results of lexical-overlapping and semantic-matching based evaluations are reported in Table 15. Similarly, the results of subjective evaluations are reported in Figure 26 and Figure 27 (see Table 16 for raw values). Some of the key findings we got are,
  • Falcon explanations got highest $R_{1}$ score for both LIAR-RAW (0.23) and RAW-FC (0.40). It means they show maximum unigram overlap with the gold explanations. Similarly, Mistral and Falcon explanations got the highest $R_{L}$ score for LIAR-RAW (0.14). Mistral explanations got highest $R_{L}$ for RAW-FC (0.18) as well. It indicates that these explanations show a maximum overlap of longest common subsequences with the gold explanations. Interestingly, we see a small deviation for $R_{2}$ scores. While explanations generated by Mistral, Llama and Falcon scored high $R_{2}$ for LIAR-RAW (0.06- 0.07) dataset, Gemma scored highest (0.20) for RAW-FC dataset. It means their explanations show maximum bigram overlap with the gold explanations.
  • As evaluating VLLMs, four out of five models found that Llama-generated explanations are better in four out of five subjective dimensions i.e. (informativeness: 4.73 - 4.78, readability: 4.51 - 4.91, objectivity: 4.17 - 4.60 and Logicality: 4.26 - 4.68) for the LIAR-RAW dataset. However, most of them found Falcon-generated explanations to be more accurate (3.55 - 3.96). Similarly, three out of five evaluating VLLMs found that Llama-generated explanations are better for all five dimensions (informativeness: 4.48 - 4.92, accuracy: 4.10 - 4.14, readability: 4.43 - 4.82, objectivity: 4.28 - 4.47 and Logicality: 4.33 - 4.52). This observations correlate with the veracity prediction results, as RoBERTa and XLNet gave the highest $MP$, MRMR and $MF1$ when trained with Llama-generated explanations.
4.3 Observations from ablation study:
In this section, we reported our observations from the ablation study. We took the best-performing models (XLNet-LLlama for LIAR-RAW and RoBERTa-LLlama for RAW-FC) for both datasets and removed individual justification (supporting and refuting) components during training to gauge their impact. Results we got for different training scenarios are reported in Table 14. Apart from that, we also segmented the test set into six parts based on the number of evidence pieces each claim has, and calculated their performances. The segment-wise performance scores are reported in Table 17 and Table 18, respectively. We observed the following,
  • Upon removing the supporting justifications from training input (see (‘w/o Supp. just.’ in Table 14), we found that $MP$, MRMR and $MF1$ scores dropped by (i) 30.90%, 29.63% and 31.48%, respectively, for LIAR-RAW (ii) 12.50%, 11.36% and 12.50%, respectively, for RAW-FC. Similarly, when we removed the refuting justifications (see (‘w/o ref. just.’ in Table 14), we saw a performance drop as well. $MP$, MRMR and $MF1$ scores dropped by (i) 10.90%, 7.41% and 9.26% respectively for LIAR-RAW (ii) 9.09%, 9.09% and 9.09% respectively for RAW-FC. Finally, when we removed both justifications, the models performed worse. The $MP$, MRMR and $MF1$ scores dropped by (i) 52.73%, 51.85% and 55.56% respectively for LIAR-RAW (ii) 47.73%, 47.73% and 47.73% respectively for RAW-FC. Removing the supporting justifications had a more adversarial impact than removing refuting justifications, and removing both had the highest adversarial impact.
  • While investigating the behaviour with varying numbers of evidence, we observed the following. For LIAR-RAW, performance of veracity predictor peaked with ‘6–20’ evidences ($MP$: 0.62, $MR$: 0.60, $MF1$: 0.61). It showed significant sensitivity to increasing number of evidences with $MF1$ dropping $\textbf{29.50}\%$ from ‘6–20’ evidences (MF1: 0.61) to ‘>50’ evidences (MF1: 0.43). However, for RAW-FC, the highest $MF1$ were observed with ‘11–20’ evidences (MP: 0.93, MR: 0.95, MF1: 0.94). It showed more robustness with $MF1$ decreasing slightly $\textbf{7.45}\%$ from ‘11–20’ evidences (MF1: 0.94) to ‘>50’ evidences (MF1: 0.87).
4.4 Linguistic insights:
In this section, we have reported how our best-performing model (XLNet-LLlama for LIAR-RAW and RoBERTa-LLlama for RAW-FC) is giving attention to different words and phrases while predicting veracity. We presented some of the samples, their gold labels, predicted labels, support justifications, and refute justifications in Figure 28, Figure 29, Figure 30, Figure 31, Figure 32, Figure 33, Figure 34, Figure 35, and Figure 36. To visualise the attention, we highlighted the top 25% words which received higher attention scores from the respective language models in TBE-3. Here, the higher intensity of blue colour indicates a higher attention score. Some of the observations are,
  • In Figure 28 where the gold and predicted label is ‘true’, we observed high attention scores on the words like ‘supported by multiple sources’, ‘evidence’, ‘supports’ in the support justification part. However, attention scores are more scattered in refute justification. Also, in contrast to supporting justification, refuting justification doesn’t consist of any phrases which refute the claim with higher confidence.
  • In Figure 29, where the label is ‘mostly-true’, attention scores were high on quantitative results such as percentage scores in the supportive justification. But in refutive justification, LLM focused on counter-evidence with less intensity of attention.
  • In Figure 30, where the label is ‘half-true’, we observed the LLM gave attention to some similar keywords in both support and refute justification. For example, ‘Gov’, ‘education’, etc. are prominent ones. Also, we observed more attention towards phrases like ‘appears that the claim is accurate’ in support justification and ‘inconsistencies undermine the validity of the claim’ in refute justification. It justifies the predicted label as half-true.
  • In Figure 31 where the label is ‘barely-true’, the VLLM generated justification focused on references to Obamacare and how many health plans were canceled nationwide. It noted Florida may have been heavily affected because of its size. But it also noticed that the exact number (300,000) is not confirmed. Subsequently, the LLM also gave more attention to keywords like ‘Obamacare’ due to which the LLM may have understood the claim might be true. Similarly, in the refute justification, VLLM confidently contradicted the claim by saying that ‘customers were not immediately dropped’. As a result, the LLM extends more attention to ‘false’ and ‘coverage’, marking the reasons to doubt the claim. Thus, the LLM figured out that the claim is not entirely false, but also not fully accurate and labelled it as ‘barely-true’.
  • In Figure 32 where the label is ‘false’, the LLM assigned higher attention scores to economic indicators, due to the absence of reasons to support the claim. Whereas the refuting justification consisted of statements disproving the claim and LLM showed more attention toward terms like ‘layoffs’, ‘assertion’, and ‘deaths’, which provide counter-evidence to refute the claim.
  • In Figure 33 where the label is ‘pants-fire’, the support justification is not useful as it is filled with repetition and irrelevant chat-bot fillers like ‘anything else I can help you with’, ‘let me know I can assist you further’, etc. Thus, we observed more attention scores on non-useful stopwords. While in the refuting justification, the VLLM argued with authority that the claim is misleading and lacks merit. We observed high attention scores on the phrase ‘the claim is unfounded and lacks merit’ which aligns with the predicted label ‘pants-fire’.
  • In Figure 34, where the label is ‘true’, the support justification generated by VLLM emphasised on the evidences like Fintan O’s (well-known columnist) writing and reputation of ‘Irish Times’. The LLM also assigned higher attention scores to these details, which supports the claim with authority. However, refute justification seems to be weaker in disproving the claim due to the lack of refuting evidence.
  • In Figure 35, where the label is ‘half’ (denoting half-true), the LLM put high attention on keywords like ‘evidence’, ‘statement is true’, ‘credibility’ and ‘support’ around factual components justifying truthfulness in the support justification. While in the refute justification, we observed higher attention scores on keywords like ‘false’, and phrases like ‘disputes the claim’ and ’In conclusion, the evidence suggests that the claim is false’. That is how the LLM evaluated the conflicting narratives and reached to a ‘half’ label.
  • In Figure 36, where the label is ‘false’, VLLM did not generate any factual justification for supporting the claim. Subsequently, the LLM could not find any solid fact or figure to put more attention on it. On the other hand, in the refute justification, VLLM justified with facts like event being from 2011, not during COVID. Also, the horses were rehomed well before the claim was made. Thus, the LLM could extend higher attention scores to phrases like ‘claim is not supported by facts’ and ‘outdated story’.
@@ 5 Conclusion:
We drew the following conclusions from our experiments and analysis,
@@
  • The role of VLLM entailed justification as a second step in TBE-3 is justified in the ablation study, where we observed that removing supporting and refuting justification adversarially affected the scores.
  • In the linguistic analysis of model explanations, we found that LLM could attend to supporting and refuting keywords and factual pieces of information for samples labelled with ‘true’ and ‘false’ labels. However, for samples with other veracity labels, attention seem to be scattered.
@@ 6 Limitations:
  • We restricted ourselves to evaluating our approach on the LIAR-RAW and RAW-FC datasets. In the future, one can modify our approach to check its performance on a variety of fact-checking datasets like FEVERThorne et al. (2018) series, fact-check-bench Wang et al. (2024c), etc.
@@ Appendix A Related works:
The NLP community has long been focusing on the core and peripheral tasks of automated fact-checking. Several datasets and methodologies have been proposed in the past, spanning different domains, languages, and task frameworks. In the following, we have reported the past literature for (i) monolingual fact-checking, (ii) multi-lingual and multi-modal fact-checking, (iii) explainable fact-checking, and (iv) other associated tasks in this area. Apart from that, we have also reported how our work fills the current research gap.
@@ Appendix B Additional details on datasets:
Figure 3: Illustration of number of claims v/s number of evidences for LIAR-RAW dataset
@@
Table 6: A sample from the LIAR-RAW dataset. It contains five key fields: (i) an identifier ‘Event_id’, (ii) the ‘claim’ to be fact-checked, (iii) the ground-truth ‘Label’, (iv) an explanation (‘Explain’) that clarifies its ground truth, and (v) the evidence sentences (E1–E4) extracted from the underlying fact-checking article.
@@
Table 7: A sample from the RAW-FC dataset. It contains five key fields: (i) an identifier ‘Event_id’, (ii) the ‘claim’ to be fact-checked, (iii) the ground-truth ‘Label’, (iv) an explanation (‘Explain’) that clarifies its ground truth, and (v) the evidence sentences (E1–E4) extracted from the underlying fact-checking article.
Figure 4: Illustration of number of claims v/s number of evidences for RAW-FC dataset
@@ C.4 IBE-4: Prompting based on entailment:
In the last inference-based experiment, we tried to answer "Can prompting with VLLM generated entailed-justifications enhance language models’ ability to predict claim veracity?" To conduct this experiment, we have followed a three-step approach similar to TBE-3 as described in Section 3.1.3. While the initial two steps, i.e. (i) to classify the evidence sentences as supporting or refuting a given claim and (ii) generating the justifications based on classified evidence sentences, are exactly the same, in the last step, instead of training, we prompted the VLLMs to generate the veracity. In past, researchers have prompted the VLLMs to find entailment for tasks like claim matching Choi and Ferrara (2024b) and counterfactual generation Dai et al. (2022). However, to the best of our knowledge, we are the first to (i) apply it to classify each evidence into supporting or refuting categories, and (ii) generate entailed justifications and use them for fact verification. The detailed approach is illustrated in Fig. 2 (c). Some of the prompt samples are show cased in the Figure 13, Figure 14, Figure 15, Figure 16 and 17.
@@ C.6.4 Evaluation metrics:
Since we are working in a multi-class classification framework for veracity prediction, we used standard metrics such as macro-precision ($MP$), macro-recall ($MR$), and macro-f1 ($MF1$) to evaluate our models. To evaluate model explainability, we first concatenated the supporting and refuting entailed justifications (generated as part of TBE-3) and considered them as model explanations. The evaluation was done with two types of strategies: (i) checking lexical overlap and semantic matching, and (ii) doing subjective evaluation by VLLMs. To check lexical overlap, we used several standard evaluation metrics such as ROUGE-1 ($R_{1}$), ROUGE-2 ($R_{2}$), ROUGE-L ($R_{L}$) Lin (2004) and BLEU Papineni et al. (2002). While $R_{1}$ and $R_{2}$ measure the overlap of unigrams and bigrams between predicted and gold explanations, $R_{L}$ measures the longest common subsequence. BLEU, on the other hand, measures the precision of matching n-grams between predicted and gold explanations. To measure the semantic matching between predicted and gold explanations, we deployed BERT score Zhang et al. (2019). It generates contextual embeddings of predicted and gold explanations and calculates cosine similarity between them. To do the subjective evaluation, we prompted the considered VLLMs (a.k.a evaluating VLLMs) to asses the model explanations generated by each VLLM (generating VLLM). The VLLMs were asked to asses across five dimensions (i) informativeness, (ii) logicality, (iii) objectivity, (iv) readability and (v) accuracy Zheng et al. (2025). The prompt template used for the assessment is presented in Figure 25.
@@ Appendix D Additional Results and Discussion:
D.1 Additional observations for macro precision in veracity prediction:
We reported the macro precision ($MP$) of various models under TBE-1, TBE-2TBE-3, and IBEs in Table 11, Table 12, and Table 13, respectively. Figure 23 illustrates a comparison of the best-performing models across them. Our observations are as follows:
  • In the case of LIAR-RAW, none of the macro-precision scores presented by IBE models surpassed the baseline performance. In contrast, for RAW-FC, the performance of models in IBE-1 (Qwen, $MP$: 0.61), and IBE-2 (Llama, $MP$: 0.62) was on equal with the highest performing baseline ($MP$: 0.62), but did not exceed it.
@@ D.1 Additional observations for macro precision in veracity prediction:
  • In TBE-1, for LIAR-RAW, Mistral model with LoRA adapter reported the highest $MP$, but did not able to surpass baseline performance. However, for RAW-FC, we found that many models, such as (i) Mistral ($MP$: 0.69, $\sim\textbf{11.29}\%\uparrow$), Llama ($MP$: 0.68, $\sim\textbf{9.68}\%\uparrow$), and Qwen ($MP$: 0.67, $\sim\textbf{8.06}\%\uparrow$) trained with LoRA, and (ii) Llama ($MP$: 0.67, $\sim\textbf{8.06}\%\uparrow$) and Qwen ($MP$: 0.70, $\sim\textbf{12.90}\%\uparrow$) trained with LoRA+, outperform the best baseline performance ($MP$: 0.62).
@@
  • In TBE-2, for LIAR-RAW, the highest reported $MP$ (0.36) does not surpass the best baseline performance ($MP$: 0.47). However, for the RAW-FC dataset, we observed that many models, such as (i) XLNet fine-tuned on Llama understandings ($\sim\textbf{1.61}\%\uparrow$), and (ii) Mistral ($\sim\textbf{6.45}\%\uparrow$) and Llama ($\sim\textbf{17.74}\%\uparrow$) trained with their respective understandings and LoRA+ adapters outperformed the best reported baseline MPMP (MPMP: 0.62).